briyamilet1:

Describe how to transform the quantity of the fifth root of x to the seventh power, to the third power into an expression with a rational exponent. Make sure you respond with complete sentences.

briyamilet1:

help anyone? xD

corey777:

Fifth root of x is x^1/5 as an exponential expression. . This to the 7th power is x ^(1/5 * 7) which is x^(7/5). Bringing this to the third power is x^ (21/5) which is your answer.

justjm:

Use this rule for exponents: \(\sqrt{x^{a}}=x^{\frac{a}{b}}\)

briyamilet1:

.-. ok... thanks ig? im a lil slow but i'll figure it out xD

justjm:

\(\sqrt{x^{a}}=x^{\frac{a}{b}}\)

justjm:

wait is it me or is the latex not working

Razor:

1 attachment
Razor:

He means that

briyamilet1:

ok... i think i get it now... ty :P

justjm:

justjm:

yeah thx razor :)

briyamilet1:

:P ty peopleee